MATH 367, spg 98, CHAPTER 2 Set II, Jan 30

Multinomials, Distributions of Balls in Urns, miscelaneous.

1. How many non–negative integer valued solutions of x_{1} +x_{2} +x_{3} = 10 are possible?

2. 8 new teachers are to be divided among 4 schools.

(a) How many divisions are possible?

(b) How many if each school must receive at least one teacher?

(c) How many if each school must receive 2 teachers?

For next class, Mon Feb 2: read chapter 2 and do problems 19, 20 21, 22, 23, 26, 28, 33, 34, 36, 38, 39, 40, 42, 43, 47, 55